The demand and total cost functions of a good are respectively and



Find expressions for TR, (profit) , MR, and MC in terms of Q.

Solve the equation


and hence determine the value of Q which maximizes profit.

Verify that, at the point of maximum profit, MR=MC.


1
Expert's answer
2021-05-25T16:55:59-0400

complete question

The demand and total cost functions of a good are

4P+Q-16=0

TC=4+2Q-3Q2/10+Q3/20

1:Find expressions for TR, (profit) , MR, and MC in terms of Q.

2:Solve the equation   dr/dQ=0 and hence determine the value of Q which maximizes profit.

3:Verify that, at the point of maximum profit, MR=MC.


solution

(1)"TR=P\times Q"

"=[4-\frac{Q}{4}]\times Q"

"=4Q-\frac{Q^{2}}{4}"


"MR=\frac{\delta TR}{\delta Q}"

"MR=\frac{\delta 4Q-\frac{Q^{2}}{4}}{\delta Q}"

"MR=4-\frac{Q}{2}"


"(profit)\pi=TR-TC"

"=[4Q-\frac{Q^{2}}{4}]-[4+2Q-\frac{3Q^{2}}{10}+\frac{Q^{3}}{20}]"

"=2Q-\frac{Q^{2}}{4}+\frac{3Q^{2}}{10}-\frac{Q^{3}}{20}-4"

"=2Q-\frac{Q^{2}}{4}+\frac{5Q^{2}+6Q^{2}}{20}-\frac{Q^{3}}{20}-4"

"\pi=2Q+\frac{Q^{2}}{20}-\frac{Q^{3}}{20}-4"


"MC=\frac{\partial TC}{\partial Q}"

"=\partial\div \partial Q[4+2Q-\frac{3Q^{2}}{10}+\frac{Q^{3}}{20}]"

"MC=2-\frac{6Q}{10}+\frac{3Q^{2}}{20}"


(2) "\frac{\partial \pi}{\partial Q}=\pi^{"}=0"

"=\partial\div \partial Q[2Q+\frac{Q^{2}}{20}-\frac{Q^{3}}{20}-4]=0"

"2+\frac{2Q}{20}-\frac{3Q^{2}}{20}=0"

"40+2Q-3Q^{2}=0"

"3Q^{2}-2Q-40=0"

"3Q^{2}-12Q+10Q-40=0"

"3Q(Q-4)+10(Q-4)=0"

"(3Q+10)(Q-4)=0"

"Q=\frac {-10}{3}"

"Q=4"

Since Q could never be negative therefore,-10/3 is rejected.

Hence=4 maximizes profit.


(3)

"MR=4-\frac{Q}{2}"

at Q=4 "MR=4-\frac{1}{2}(4)=2"

"MC=2-\frac{6Q}{10}+\frac{3Q^{2}}{20}"

at Q=4 "MC=2-\frac{6\times4}{10}+\frac{3\times 4^{2}}{20}=2"

at Q=4

"MR=MC=2"